# Adaptive Inference-Time Scaling via Cyclic Diffusion Search

### v3

@@ 2 Preliminaries
Figure 1: Overview of our method. (1) Start with $N$ particles, use jumpy denoising $x_{T}\rightarrow x_{T-j}\rightarrow\cdot\cdot\cdot\rightarrow x_{0}$. (2) Replicate each particles $J$ times and send each particles to multiple noise levels (AEEB). (3) Re-denoise particles. (4) Select Top $K$ particles. (5) We repeat steps (2)–(4) until the adaptive terminal condition is satisfied (CDS). The number of cycles executed in this way defines the Adaptive Thinking Time (ATT).
@@ 3.2 Automatic Exploration-Exploitation Balancing
2:     Initialize $N$ particles $\{x_{T}^{(i)}\}$ from ${\cal N}(0,I)$
@@ 5 Experiments
We evaluate our proposed method on a diverse suite of challenging tasks that require generating samples in sparse or previously unexplored regions at training phase: (1) toy Mixture of Gaussian(MoG), (2) point-mass maze navigation (park2024ogbench, ), (3) Sudoku puzzle completion (palm2018recurrentrelationalnetworks, ; wang2019satnetbridgingdeeplearning, ), (4) path generation on unseen pixel maze image (maze, ), and (5) molecular structure prediction (hoogeboom2022, ). These tasks present distinct challenges: MoG shows the importance of multiple go-back temperature pool as the proof-of-concept; point-mass maze requires long-horizon planning over 1000 steps; Sudoku demands logical consistency across row, column, and block constraints; path generation on unseen maze image tests generalization to novel environmental structures; and molecular structure prediction requires generating valid 3D conformations under chemical and physical constraints. Detailed task configurations are discussed in Appendix B.2.
@@ 5.1 Baselines
All baselines used the same pre-trained diffusion model per tasks, and inference-time budgets were expanded as much as possible while preserving their core operational characteristics. In BS and SoP, inference-time compute was increased by more frequent expansion and selection, while SMC scaled compute via higher resampling frequency, and BoN by increasing the number of particles $N$. For fair comparison, the total number of particles was set equal across all methods (except BoN). Additional experimental details for each tasks are provided in the Appendix B.6B.8B.10B.12.
Figure 2: Evaluated Tasks. We evaluate ABCD on one toy tasks and four challenging tasks requiring generation in sparse regions of solution space: (a) Mixture of Gaussian for Proof-of-concept that we need multiple go back temperature. (b) Sudoku puzzle completion demanding logical consistency across multiple constraints simultaneously; (c) Pixel Maze Path Finding testing generalization to novel environmental structures not encountered during training; (d) Molecular structure prediction requiring physically and chemically valid 3D conformations. (e) OGBench PointMaze navigation requiring complex planning over 1000+ steps;
@@ 5.2 Mixture of Gaussian
Setup. We construct two toy datasets to illustrate the need for multiple go-back noise levels and terminal condition during inference refinement (Figure 2(a)). Dataset 1 has modes clustered locally, so small-step refinement suffices; Dataset 2’s modes lie in distant regions, requiring occasional large “jumps” to escape poor initial prediction. We train separate diffusion models on each and condition sampling on distance to a target (red star). Full details appear in Appendix B.4.
Result. We evaluate ABCD against four fixed single go-back temperatures (GB 20, 40, 60, 80) under two inference regimes: a fixed 25-cycle budget and an open-ended budget of up to 100 cycles with early stopping upon goal attainment. As shown in comparison with fixed 25 cycles (Table 2): only ABCD attains 100 % success and minimal final distance on both datasets. Fixed small go-back steps work on Dataset 1 but cannot escape poor initializations in Dataset 2, whereas fixed large steps succeed escaping from poor initializations on Dataset 2 but over-cycle on Dataset 1 and 2 (e.g., reaching the goal at cycle 20 but diverging by cycle 25 after losing useful information). As shown in comparison with up to 100 cycles with early stopping (Table  2): some fixed levels solve one dataset, but only ABCD consistently achieves both 100 % success with comparable average cycles. These findings underscore that exploration (coarse jumps) and exploitation (fine steps) require different go-back levels—a balance that ABCD’s adaptive termination naturally realizes.
@@ 5.3 Sudoku puzzle completion
Setup. We evaluated ABCD against baseline methods on the Sudoku puzzle completion task (palm2018recurrentrelationalnetworks, ; wang2019satnetbridgingdeeplearning, ), a canonical logical-reasoning benchmark. In Sudoku, a 9×9 grid must be filled under row, column, and 3×3-block constraints, with each new grid filled based on the current grid state. An adaptive search strategy—broadly exploring when few clues are given and then locally refining once most entries are correct—proves particularly effective. We trained all models on puzzles with 31–42 given digits wang2019satnetbridgingdeeplearning and tested them on more challenging instances containing only 17–28 givens palm2018recurrentrelationalnetworks , thereby measuring each method’s ability to produce logically consistent solutions. See Appendix B.8 for additional experimental details.
@@ 5.4 Path generation on unseen pixel maze image
Setup. We evaluated ABCD’s adaptability on a path-generation task using mazes of previously unseen sizes (maze, ). Every method leveraged the same diffusion model pretrained on small mazes (sizes 4, 5, and 6) and was then tested on larger, structurally distinct mazes zhang2025tscendtesttimescalablemctsenhanced . This setting rigorously probes each algorithm’s capacity to generalize to novel spatial configurations at inference time. Additional experimental details are provided in Appendix B.10.
Figure 5: Pixel Maze path finding results on multiple maze size. Success rate comparison across maze sizes in the Pixel Maze task. All settings involve OOD evaluation, where models are trained on smaller mazes (sizes 4$\sim$6) and tested on larger ones. As maze size increases, the search space grows and the performance gap between other methods widens, highlighting the importance of adaptive inference-time exploration in complex scenarios.
Result. As shown in Figure 4(Middle), ABCD reaches near-perfect success rates significantly faster than all baselines, evidencing a superior time–accuracy trade-off. In contrast, most competing methods—including Base Diffusion, BoN, BS, and SMC—never exceed a 60 % success rate even under generous computational budgets. ABCD, by comparison, attains 100 % success in substantially less time. Figure 5 further examines performance across maze sizes 11–15 in the Pixel Maze task. Here again, ABCD consistently dominates both success rate and inference time at every size, and the margin over baselines widens as the search space grows. These findings underscore the value of adaptive, instance-specific exploration strategies especially in complex search demanding tasks.
@@ 5.7 Do we need to care about go-back temperature level?
Instance-Level optimal temperature. Figure 7(a–b) plots cycle-wise accuracy for two Sudoku instances (14 vs. 44 givens) under fixed go-back noise levels. Certain noise settings plateau after only a few cycles and the plateau point differs between puzzles, indicating that each instance possesses its own optimal noise level. (see Appendix C.1).
Temperature selection per cycle. Figure 8 tracks the provenance of selected particles source across cycles for three instance types in Sudoku (17, 19, and 32 givens). Harder puzzles continue sampling from a broad range of noise levels, while easier ones converge rapidly to low-variance perturbations. Early iterations emphasize high-variance (global) moves; later iterations favor low-variance (local) refinements. These patterns demonstrate ABCD’s dynamic, per-instance generative process through adaptive exploration vs. exploitation.
@@ 5.8 Do we need adaptive terminal condition?
Efficiency of adaptive terminal condition. An ablation study on Pixel maze size 15 (Figure 7(c) compares fixed cycle length inference (blue) to our adaptive terminal criterion (red) with the same temperature pool. The adaptive policy stops early on simple puzzles—avoiding wasted compute—yet allocates extra iterations to hard ones, yielding consistently superior time–success trade-offs.
ABCD automatically assigns proper thinking time. Figure 7(d) demonstrates that ABCD allocates more thinking cycles to harder puzzles (fewer givens) and terminates earlier on simpler ones, validating that our adaptive terminal criterion scales compute with problem difficulty. Furthermore, the stop-flag hyperparameter $\kappa$ directly controls exploration depth: larger $\kappa$ values enforce additional cycles before stopping, which boosts success on challenging instances. Identical trends hold on the Pixel Maze task. (see Appendix B.11).
@@ B.2 Task Detail
We introduce the setting of experiments and diffusion models we used. For Mixture of Gaussian, we constructed the two datasets generated from the mixture of gaussians. Dataset 1 is generated from 1 region of Gaussians and Dataset 2 is generated from 2 region of the islands as you can see on the Figure 2(a). After pre-training two diffusion models on each dataset, we guide sampling at inference time by providing the distance to a target goal as a reward, encouraging generation toward the desired goal state. For OGBench Maze, we adopted the task setups from [18], running 5 sparse reward navigation tasks in each environment, each can take hundreds of steps to accomplish. The model is trained on standard dataset collected by noisy expert policy that repeatedly reaching randomly sampled goals. At test time, the model is required to generate proper plan using guidance from value function. For Sudoku, the basic setting is adopted from IRED [7], where the diffusion model is trained on SAT-Net dataset with 31 to 42 provided samples [27] and tested on harder RRN dataset [17] with 17 to 28 provided samples. For Pixel Maze, the basic setting is adopted from T-SCEND [31], where the dataset are generated by maze-dataset [11], and the diffusion model is trained with Maze sizes of 4 $\times$ 4 to 6 $\times$ 6 and tested with size of $11\times 1$ 11 to $15\times 15$. For Molecule Generation, the basic setting is adopted from EDM [10]. The diffusion model is trained to generate 3D molecular geometries using the QM9 dataset [19], which contains approximately 130k small molecules, each with up to 9 heavy atoms. We utilized the standard splits of this dataset for training, validation, and testing.
@@
Verifiers. For OGBench Maze, we employ a soft verifier that scores each generated plan based on the proportion of collision-free states throughout the trajectory and the proximity to the goal at the end of the trajectory. For Sudoku, we use the ratio of correct answer that matches the ground truth label. For Pixel Maze, we use weighted sum of Precision, Recall, F1 score with rate of success score which indicates whether the path reaches the goal without passing the wall. For Molecule Generation, we use the Negative Log-Likelihood (NLL) to evaluate the generated molecular structures.
@@
Evaluation metrics. For OGBench Maze, we used the success rate (whether the planned trajectory reaches to the goal or not) of the actual rollouts from the generated trajectory plans. For Sudoku, we use the ratio of correct answer that matches the ground truth label. For Pixel Maze, we use success score (0 or 1) which indicates whether the path reaches the goal without passing the wall. For Molecule Generation, we use molecule stability (0 or 1), the proportion of generated molecules for which all atoms are stable.
@@ B.5 MoG Result Detail
After leveraging the large go-back to traverse between regions, switching to small go-back becomes advantageous for shifting from exploration to exploitation. This transition enables fine-grained goal-oriented refinement, as demonstrated in Figure 10. All instances begin from an initial incorrect prediction, requiring additional correction during inference. go-back $=20$ fails to escape the initial cluster of modes, whereas go-back $=80$ enables transitions across regions but refines inefficiently. Our method combines coarse transitions to explore distant regions and fine adjustments for local refinement, enabling faster and more effective inference time adjustment.
@@ B.9 Sudoku Result
Inference Method Add. Compute Accuracy Inference wall clock time (sec.)
Base Diffusion No add 0.35 ± 0.1 0.28
N = 40 0.55 ± 0.49 0.29
BoN N = 480 0.63 ± 0.48 0.68
N = 960 0.66 ± 0.09 1.32
$p=5$ 0.64 ± 0.1 0.3
SMC $p=2$ 0.70 ± 0.1 0.31
$p=1$ 0.74 ± 0.1 0.32
$p=5$ 0.65 ± 0.46 0.32
BS $p=2$ 0.75 ± 0.48 0.40
$p=1$ 0.84 ± 0.08 0.54
$\Delta f=40$, $\Delta b=20$ 0.61 ± 0.09 0.41
$\Delta f=20$, $\Delta b=10$ 0.67 ± 0.1 0.46
$\Delta f=10$, $\Delta b=5$ 0.76 ± 0.1 0.51
$\Delta f=6$, $\Delta b=3$ 0.83 ± 0.11 0.52
SoP $\Delta f=4$, $\Delta b=2$ 0.88 ± 0.09 0.55
$\Delta f=2$, $\Delta b=1$ 0.95 ± 0.06 0.65
$\Delta f=4$, $\Delta b=3$ 0.955 ± 0.06 0.99
$\Delta f=5$, $\Delta b=4$ 0.958 ± 0.06 1.14
$\Delta f=6$, $\Delta b=5$ 0.957 ± 0.06 1.33
percentage = 0.6, $\kappa$ = 1 0.82 ± 0.14 0.24
percentage = 0.8, $\kappa$ = 1 0.90 ± 0.1 0.35
percentage = 1.0, $\kappa$ = 1 0.95 ± 0.06 0.45
percentage = 1.0, $\kappa$ = 2 0.97 ± 0.05 0.49
ABCD percentage = 1.0, $\kappa$ = 3 0.979 ± 0.04 0.55
percentage = 1.0, $\kappa$ = 5 0.986 ± 0.03 0.60
percentage = 1.0, $\kappa$ = 10 0.99 ± 0.01 0.75
percentage = 1.0, $\kappa$ = 15 0.995 ± 0.01 0.85
percentage = 1.0, $\kappa$ = 20 0.996 ± 0.01 0.92

### v4

@@ 2 Preliminaries
Figure 1: Overview of our method. (1) Start with NN particles, use jumpy denoising $x_{T}\rightarrow x_{T-j}\rightarrow\cdot\cdot\cdot\rightarrow x_{0}$. (2) Replicate each particles $J$ times and send each particles to multiple noise levels (AEEB). (3) Re-denoise particles. (4) Select Top $K$ particles. (5) We repeat steps (2)–(4) until the adaptive terminal condition is satisfied (CDS). The number of cycles executed in this way defines the Adaptive Thinking Time (ATT).
@@ 3.2 Automatic Exploration-Exploitation Balancing
2:  Initialize NN particles $\{x_{T}^{(i)}\}$ from ${\cal N}(0,I)$
@@ 5 Experiments
We evaluate our proposed method on a diverse suite of challenging tasks that require generating samples in sparse or previously unexplored regions at training phase: (1) toy Mixture of Gaussian(MoG), (2) point-mass maze navigation (park2024ogbench, ), (3) Sudoku puzzle completion (palm2018recurrentrelationalnetworks, ; wang2019satnetbridgingdeeplearning, ), (4) path generation on unseen pixel maze image (maze, ), (5) molecular structure prediction (hoogeboom2022, ), and (6) text-to-image generation. These tasks present distinct challenges: MoG shows the importance of multiple go-back temperature pool as the proof-of-concept; point-mass maze requires long-horizon planning over 1000 steps; Sudoku demands logical consistency across row, column, and block constraints; path generation on unseen maze image tests generalization to novel environmental structures; molecular structure prediction requires generating valid 3D conformations under chemical and physical constraints; and image generation illustrates that our approach also operates effectively in the high-dimensional setting, demonstrating its scalability beyond structured low-dimensional tasks. Detailed task configurations are discussed in Appendix B.2.
@@ 5.1 Baselines
All baselines used the same pre-trained diffusion model per tasks, and inference-time budgets were expanded as much as possible while preserving their core operational characteristics. In BS and SoP, inference-time compute was increased by more frequent expansion and selection, while SMC scaled compute via higher resampling frequency, and BoN by increasing the number of particles $N$. For fair comparison, the total number of particles was set equal across all methods (except BoN). While we primarily report wall-clock time in the main paper to analyze inference-time scaling behavior, we additionally provide a analysis of the Number of Function Evaluations (NFEs) in Appendix B.16. Additional experimental details for each tasks are provided in the Appendix B.6B.8B.10B.12.
Figure 2: Evaluated Tasks. (a) Mixture of Gaussian for Proof-of-concept that we need multiple go back temperature. (b) Sudoku puzzle completion demanding logical consistency across multiple constraints simultaneously; (c) Pixel Maze Path Finding testing generalization to novel environmental structures not encountered during training; (d) Molecular structure prediction requiring physically and chemically valid 3D conformations. (e) OGBench PointMaze navigation requiring complex planning over 1000+ steps (f) Text-to-image generation requiring complex, high-dimensional outputs that are well aligned with the given textual descriptions;
@@ 5.2 Mixture of Gaussian
Setup. We construct two toy datasets to illustrate the need for multiple go-back noise levels and terminal condition during inference refinement (Figure 2(a)). Dataset 1 has modes clustered locally, so small‐step refinement suffices; Dataset 2’s modes lie in distant regions, requiring occasional large “jumps” to escape poor initial prediction. We train separate diffusion models on each and condition sampling on distance to a target (red star). Full details appear in Appendix B.4.
Result. We evaluate ABCD against four fixed single go‐back temperatures (GB 20, 40, 60, 80) under two inference regimes: a fixed 25-cycle budget and an open-ended budget of up to 100 cycles with early stopping upon goal attainment. As shown in comparison with fixed 25 cycles (Table 2): only ABCD attains 100 % success and minimal final distance on both datasets. Fixed small go-back steps work on Dataset 1 but cannot escape poor initializations in Dataset 2, whereas fixed large steps succeed escaping from poor initializations on Dataset 2 but over‐cycle on Dataset 1 and 2 (e.g., reaching the goal at cycle 20 but diverging by cycle 25 after losing useful information). As shown in comparison with up to 100 cycles with early stopping (Table  2): some fixed levels solve one dataset, but only ABCD consistently achieves both 100 % success with comparable average cycles. These findings underscore that exploration (coarse jumps) and exploitation (fine steps) require different go-back levels—a balance that ABCD’s adaptive termination naturally realizes.
@@ 5.3 Sudoku puzzle completion
Setup. We evaluated ABCD against baseline methods on the Sudoku puzzle completion task (palm2018recurrentrelationalnetworks, ; wang2019satnetbridgingdeeplearning, ), a canonical logical‐reasoning benchmark. In Sudoku, a 9×9 grid must be filled under row, column, and 3×3‐block constraints, with each new grid filled based on the current grid state. An adaptive search strategy—broadly exploring when few clues are given and then locally refining once most entries are correct—proves particularly effective. We trained all models on puzzles with 31–42 given digits wang2019satnetbridgingdeeplearning and tested them on more challenging instances containing only 17–28 givens palm2018recurrentrelationalnetworks , thereby measuring each method’s ability to produce logically consistent solutions. See Appendix B.8 for additional experimental details.
@@ 5.4 Path generation on unseen pixel maze image
Setup. We evaluated ABCD’s adaptability on a path‐generation task using mazes of previously unseen sizes (maze, ). Every method leveraged the same diffusion model pretrained on small mazes (sizes 4, 5, and 6) and was then tested on larger, structurally distinct mazes zhang2025tscendtesttimescalablemctsenhanced . This setting rigorously probes each algorithm’s capacity to generalize to novel spatial configurations at inference time. Additional experimental details are provided in Appendix B.10.
[…]
Result. As shown in Figure 4(Middle), ABCD reaches near‐perfect success rates significantly faster than all baselines, evidencing a superior time–accuracy trade‐off. In contrast, most competing methods—including Base Diffusion, BoN, BS, and SMC—never exceed a 60 % success rate even under generous computational budgets. ABCD, by comparison, attains 100 % success in substantially less time. Figure 5 further examines performance across maze sizes 11–15 in the Pixel Maze task. Here again, ABCD consistently dominates both success rate and inference time at every size, and the margin over baselines widens as the search space grows. These findings underscore the value of adaptive, instance‐specific exploration strategies especially in complex search demanding tasks.
@@ 5.6 OGBench Point Maze
5.7 Text-to-image generation
Setup. Finally, we evaluate ABCD on the image generation tasks. We adopt Stable Diffusion v1.5, a latent diffusion model, to serve as the image prior for generating 512 × 512 samples $x\sim p_{\theta}(x|y)$, conditioned on the textual prompt y. For downstream reward function, we utilize metrics including compressibility, aesthetic evaluation (using LAION’s V2 Aesthetic Predictor schuhmann2022laion ), and human preference evaluation (HPSv2 from wu2023humanv2 ) for comprehensive evaluation. Further details regarding the experiments can be found in Appendix B.14.
Result. Figure 7 shows that ABCD consistently outperforms two strong baselines, BoN and SoP, in generating high-reward samples. It efficiently navigates the high-dimensional image space to identify high-reward samples. As computational budget increases, all methods see reward improvements, but ABCD benefits the most. For example, in all metrics, as we double or triple the inference time, BoN barely climbs, while SoP achieves moderate gains. By contrast, ABCD can hit the same compression level that SoP reaches at 272s in less than a quarter of the time. In addition, to reach a human preference score of 0.2764, SoP needs over 300s, while ABCD does it in under 81s. These highlights that its instance-specific exploration strategy effectively leverages additional computation to better align generated samples. Visualizations of the generated images are provided in Appendix B.15.1.
Figure 7: Text-to-image generation result. Average reward with respect to compressibility, aesthetic score and human preference score.
@@ 5.8 Do we need to care about go-back temperature level?
Instance-level optimal temperature. Figure 8(a–b) plots cycle‐wise accuracy for two Sudoku instances (14 vs. 44 givens) under fixed go‐back noise levels. Certain noise settings plateau after only a few cycles and the plateau point differs between puzzles, indicating that each instance possesses its own optimal noise level. (see Appendix C.1).
Temperature selection per cycle. Figure 9 tracks the provenance of selected particles source across cycles for three instance types in Sudoku (17, 19, and 32 givens). Harder puzzles continue sampling from a broad range of noise levels, while easier ones converge rapidly to low‐variance perturbations. Early iterations emphasize high‐variance (global) moves; later iterations favor low‐variance (local) refinements. These patterns demonstrate ABCD’s dynamic, per‐instance generative process through adaptive exploration vs. exploitation.
@@ 5.9 Do we need adaptive terminal condition?
Efficiency of adaptive terminal condition. An ablation study on Pixel maze size 15 (Figure 8(c) compares fixed cycle length inference (blue) to our adaptive terminal criterion (red) with the same temperature pool. The adaptive policy stops early on simple puzzles—avoiding wasted compute—yet allocates extra iterations to hard ones, yielding consistently superior time–success trade‐offs.
ABCD automatically assigns proper thinking time. Figure 8(d) demonstrates that ABCD allocates more thinking cycles to harder puzzles (fewer givens) and terminates earlier on simpler ones, validating that our adaptive terminal criterion scales compute with problem difficulty. Furthermore, the stop-flag hyperparameter $\kappa$ directly controls exploration depth: larger $\kappa$ values enforce additional cycles before stopping, which boosts success on challenging instances. Identical trends hold on the Pixel Maze task. (see Appendix B.11, C.2,C.3, E for more detailed analysis).
@@ B.2 Task detail
We introduce the setting of experiments and diffusion models we used. For Mixture of Gaussian, we constructed the two datasets generated from the mixture of gaussians. Dataset 1 is generated from 1 region of Gaussians and Dataset 2 is generated from 2 region of the islands as you can see on the Figure 2(a). After pre-training two diffusion models on each dataset, we guide sampling at inference time by providing the distance to a target goal as a reward, encouraging generation toward the desired goal state. For OGBench Maze, we adopted the task setups from [19], running 5 sparse reward navigation tasks in each environment, each can take hundreds of steps to accomplish. The model is trained on standard dataset collected by noisy expert policy that repeatedly reaching randomly sampled goals. At test time, the model is required to generate proper plan using guidance from value function. For Sudoku, the basic setting is adopted from IRED [8], where the diffusion model is trained on SAT-Net dataset with 31 to 42 provided samples [29] and tested on harder RRN dataset [18] with 17 to 28 provided samples. For Pixel Maze, the basic setting is adopted from T-SCEND [35], where the dataset are generated by maze-dataset [12], and the diffusion model is trained with Maze sizes of 4 $\times$ 4 to 6 $\times$ 6 and tested with size of $11\times 11$ to $15\times 15$. For Molecule Generation, the basic setting is adopted from EDM [11]. The diffusion model is trained to generate 3D molecular geometries using the QM9 dataset [20], which contains approximately 130k small molecules, each with up to 9 heavy atoms. We utilized the standard splits of this dataset for training, validation, and testing. For Text-to-Image generation, we use Stable Diffusion v1.5—a widely adopted text-to-image diffusion model—as our pretrained backbone.
@@
Verifiers. For OGBench Maze, we employ a soft verifier that scores each generated plan based on the proportion of collision-free states throughout the trajectory and the proximity to the goal at the end of the trajectory. For Sudoku, we use the ratio of correct answer that matches the ground truth label. For Pixel Maze, we use weighted sum of Precision, Recall, F1 score with rate of success score which indicates whether the path reaches the goal without passing the wall. For Molecule Generation, we use the Negative Log-Likelihood (NLL) to evaluate the generated molecular structures. For Text-to-Image generation, we use three verifier scores: compressibility [2], aesthetic [22], and human preference score [33].
@@
Evaluation metrics. For OGBench Maze, we used the success rate (whether the planned trajectory reaches to the goal or not) of the actual rollouts from the generated trajectory plans. For Sudoku, we use the ratio of correct answer that matches the ground truth label. For Pixel Maze, we use success score (0 or 1) which indicates whether the path reaches the goal without passing the wall. For Molecule Generation, we use molecule stability (0 or 1), the proportion of generated molecules for which all atoms are stable. For Text-to-Image generation, we use three evaluation metrics: compressibility  [2], aesthetic [22], and human preference score [33]—which correspond to specific verifiers, respectively.
@@ B.5 MoG result detail
After leveraging the large go-back to traverse between regions, switching to small go-back becomes advantageous for shifting from exploration to exploitation. This transition enables fine-grained goal-oriented refinement, as demonstrated in Figure 11. All instances begin from an initial incorrect prediction, requiring additional correction during inference. go-back =20=20 fails to escape the initial cluster of modes, whereas go-back =80=80 enables transitions across regions but refines inefficiently. Our method combines coarse transitions to explore distant regions and fine adjustments for local refinement, enabling faster and more effective inference time adjustment.
@@ B.9 Sudoku result
Inference Method Add. Compute Accuracy Clock Time (sec.) NFE
Base Diffusion No add 0.35 ± 0.1 0.28 50
N = 40 0.55 ± 0.49 0.29 2000
BoN N = 480 0.63 ± 0.48 0.68 24000
N = 960 0.66 ± 0.09 1.32 48000
$p=5$ 0.64 ± 0.1 0.3 2000
SMC $p=2$ 0.70 ± 0.1 0.31 2000
$p=1$ 0.74 ± 0.1 0.32 2000
$p=5$ 0.65 ± 0.46 0.32 900
BS $p=2$ 0.75 ± 0.48 0.40 1500
$p=1$ 0.84 ± 0.08 0.54 2460
$\Delta f=40$, $\Delta b=20$ 0.61 ± 0.09 0.41 2940
$\Delta f=20$, $\Delta b=10$ 0.67 ± 0.1 0.46 3420
$\Delta f=10$, $\Delta b=5$ 0.76 ± 0.1 0.51 3860
$\Delta f=6$, $\Delta b=3$ 0.83 ± 0.11 0.52 3940
SoP $\Delta f=4$, $\Delta b=2$ 0.88 ± 0.09 0.55 4220
$\Delta f=2$, $\Delta b=1$ 0.95 ± 0.06 0.65 4980
$\Delta f=4$, $\Delta b=3$ 0.955 ± 0.06 0.99 8260
$\Delta f=5$, $\Delta b=4$ 0.958 ± 0.06 1.14 9900
$\Delta f=6$, $\Delta b=5$ 0.957 ± 0.06 1.33 11540
percentage = 0.6, $\kappa$ = 1 0.82 ± 0.14 0.24 1059
percentage = 0.8, $\kappa$ = 1 0.90 ± 0.1 0.35 1667
percentage = 1.0, $\kappa$ = 1 0.95 ± 0.06 0.45 2296
percentage = 1.0, $\kappa$ = 2 0.97 ± 0.05 0.49 2607
ABCD percentage = 1.0, $\kappa$ = 3 0.979 ± 0.04 0.55 2827
percentage = 1.0, $\kappa$ = 5 0.986 ± 0.03 0.60 3205
percentage = 1.0, $\kappa$ = 10 0.99 ± 0.01 0.75 3942
percentage = 1.0, $\kappa$ = 15 0.995 ± 0.01 0.85 4499
percentage = 1.0, $\kappa$ = 20 0.996 ± 0.01 0.92 5013